OpenStudy (anonymous):

Two Algebra 2 questions? I give medals!

OpenStudy (anonymous):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

Solve the system by substitution. -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (anonymous):

please write steps line by line. i need to learn how to do these

OpenStudy (anonymous):

@SolomonZelman please help me I'm so uneducated:/ i have an F in algebra 2 please help me:(

OpenStudy (anonymous):

Failed my test:(